OpenStudy (sierra_tibbetts):

Maria is comparing the prices of two window cleaning companies. Company A charges $6 per window and an additional $12 as service charges. Company B charges $5 per window and an additional $15 as service charges. Part A: Write equations to represent Company A's and Company B's total charges for cleaning a certain number of windows. Define the variable used in the equations. Part B: Which company would charge less for cleaning 8 windows? Justify your answer. Part C: How much money is saved by using the services of Company B instead of Company A to clean 6 windows?

OpenStudy (cp123):

Part A: company A equation: 6x+12=T. x= number of windows, t = total. Company B equation: 5x+15=T. x= number of windows, t = total. Part B: replace x with 8 to get T, so for company A it would be 6(8)+12=60 Company B would be 5(8)+ 15 = 55, so, company B would charge less. Part C: do the same, replace x in the original equations with 6. company a would cost a total of 48 dollars, company B would cost a total of 45 dollars. there is a three dollar difference between the two of them. It is simply, really.